Project

General

Profile

NDOS Commissioning Meeting 2011/10/10

Attendance: Jon, Martin, Peter, Mark Mathis, Denis, Mark Messier, Rafael, Enrique, Leon, Susan, Minerba, Zeb, Nick

Agenda:

  1. APD Redistribution Status & Plans
  2. Beam & Detector Update
  3. DAQ Status & Plans
  4. DCS Status & Plans
  5. AOB

APD Redistribution Status & Plans

  • Jon/Mat: of the 15 redistributed APDs, we have lost 2, one of which could be directly attributed to the move. The other 13 APDs are functioning fine. We will therefore proceed with the move of more APDs. Mat and Denis are able to do about 10/day, so we will move 10 today and 10 tomorrow, and then take a break for the Collab. Meeting.
  • Mat: boroscope is out for repairs but we have a spare. The spare isn't as nice the one we used before. Will plan to instrument side of detector where it's easier to install.
  • Jon: how many open slots do we have on the side of the fiducial volume? Mat: ~15 in position two.
  • Mat: four non-cooled, recovered APDs (two coated, two non-coated) are still functioning well. We will enable cooling on these today.
  • Peter: during software updates, we blow away DCMControl. Can this package be ~stand-alone?

Beam & Detector

  • BNB is down today, there is a problem with the BNB horn.
  • NuMI beam is running fine, around 37.e12
  • Two FEBs are in temperature alarm. Martin: one was in and out of alarm during the weekend. Should be safe to raise threshold to 45 deg. C. Leon: parylene-coated APDs may require slightly higher drive current (thermal resistivity is higher)

DAQ Status & Plans

  • Gennadiy has an updated version of DDS that will address several issues, in particular some significant memory leaks. The plan is to swap out the current base release with a modified version built against the new version of DDS. However, once in place, DDS should not be restarted by the shifter, and no software updates will happen until probably next Monday (due to Collab. Mtng). We will begin these tests today.
  • Peter: dcm-3-2-1 was not producing data at various times. Andrew found the dcm to be in internal timing mode this morning, but (a) that doesn't jive with the timing markers seen during the weekend, and (b) this was after Jon did a DSO scan this morning. The timing mode is forced back into external timing mode during the "Configure Hardware" stage.

DCS Status & Plans

  • Martin: group plans to setup backups on DCS machines; should not effect anyone.

AOB

  • Susan: there was an entry in the ECL about fibers sticking out of optical connection?
  • Leon: found cleaning rag was sticking to face of optical connector while cleaning APD last week. This was in block 2, plane 21, pos. 0
  • Martin: will update IOC and conversion of HV adjust values.